# RAG-Enhanced Collaborative LLM Agents for Drug Discovery

### v2

@@ F.2 Additional External Knowledge Analysis
F.3 Additional Case Studies
In this section, we provide additional case studies to analyze the behavior CLADD. In Figure 9, we observe that all three agents consistently predict dopamine-related and serotonin-related proteins as targets. Based on the reports, Prediction Agent prioritizes these proteins over Cytochrome P450-related enzymes in their predictions. Thus, we argue that our system efficiently prioritizes relevant information based on consensus, functioning similarly to a majority voting system.
@@ Appendix G Agent Templates
In this section, we provide the templates for each agent used in Section 3.
Table 7: Prompts for Molecule Annotation Planner (Section 3.2.1).
@@
Table 8: Prompts for Knowledge Graph Planner (Section 3.2.1).
@@
Table 9: Prompts for Biology Relation Agent (Section 3.2.2).
@@
Table 10: Prompts for Drug Relation Agent (Section 3.2.2).
@@
Table 11: Prompts for Molecule Understanding Agent (Section 3.2.3).
@@
Table 12: Prompts for Prediction Agent (Section 3.2.4).

### v3

@@ F.3 Additional External Knowledge Analysis
F.4 Additional Case Studies
In this section, we provide additional details on case studies to analyze the behavior CLADD. In Figure 12, we observe that all three agents consistently predict dopamine-related and serotonin-related proteins as targets. Based on the reports, Prediction Agent prioritizes these proteins over Cytochrome P450-related enzymes in the prediction. Thus, we argue that our system can efficiently prioritize relevant information based on consensus, functioning similarly to a majority voting system.
@@ Appendix G Agent Templates
In this section, we provide the templates for each agent used in Section 3. We follow the previous work for designing the system prompt (li2024empowering).
Table 8: Prompts for Molecule Annotation Planner (Section 3.1).
@@
Table 9: Prompts for Knowledge Graph Planner (Section 3.1).
@@
Table 10: Prompts for Biology Relation Agent (Section 3.1).
@@
Table 11: Prompts for Drug Relation Agent (Section 3.1).
@@
Table 12: Prompts for Molecule Understanding Agent (Section 3.1).
@@
Table 13: Prompts for Prediction Agent (Section 3.1).